# Are Biological Systems More Intelligent Than Artificial Intelligence?

### v5

@@ 2.4 Reasons to Suspect Delegation Facilitates Adaptation
I will now give a brief explanation of why I believe delegation of control to lower levels of abstraction is relevant to adaptation.
Throughout this subsection I use examples (e.g. distributed drone swarms, mission command, and economies) as intuition-building thought experiments. They are not intended as comparative empirical evidence for the formal claim, which is established later in Theorem 5.1. All else being equal the lower the level of abstraction at which the system can be controlled, the more it can be specialised. For example a program written in C can be far more efficient than a program written in Python [oden2020]. This is because the Python program has to be interpreted by the Python interpreter, which is a program written in C. By programming in C I simply cut out an abstraction layer. This gives the programmer more leeway to optimise the system, or make mistakes. Likewise another way to make computers more efficient is to use a FPGA to write a program at the hardware level, instead of compiling that program into machine code and running it on a one-size-fits-all instruction set architecture like x86 [nguyen2020].
With each successive abstraction layer introduced, one is further removed from the ground truth to which one must adapt. For example software is nothing more than a state of hardware [^19] That software is a state of hardware has been addressed at length elsewhere [bennett2024a, bennett2025thesis], beyond the scope of this paper.. It is a label I ascribe to the behaviour of hardware. It is a useful simplification that lets us consider what I want hardware to do in the abstract, but working at such a high level of abstraction can cause problems. Proposals for software ‘intelligence’ have been undermined by the fact that any claims regarding the behaviour of software are contingent upon hardware. The software doesn’t interact with the environment, the hardware does. This is because the ability to adapt is related to perceived complexity, which is entirely determined by the abstraction layer through which an agent interacts with its environment [bennett2024b]. Consequently an ‘objective’ measure of artificial intelligence [leike2015, bennett2024a, bennett2025b] must be ‘enactive’ [thompson2007] in that it hinges not just software but hardware and the local environment as well. To understand the difference between biological systems and artificial intelligence, I must account for the level of abstraction at which an adaptive process takes place.
@@
In distributed computational systems, like a drone swarm or network, delegating control can facilitate what is called ‘graceful degradation’ [kuniavsky2010]. For example, to control a drone swarm I could single out one drone to be the ‘leader’, and have it run a program that co-ordinates the other drones. However that would make the swarm vulnerable. If the leader drone were damaged, the swarm would de-cohere. It is a single point of failure. In contrast if I can get the swarm to co-ordinate through entirely decentralised control, delegating to the level of individual drones, the swarm will remain coherent regardless of how many drones are disabled [^20] This sort of decentralised swarm control can also be observed in nature, such as with flocks of starlings which fly in complex formations [cavagna2010].. In other words, the swarm ‘adapts’ to adverse conditions.
@@
The advantages of delegating control are visible in human organisations as well. In military doctrine, “mission command” delegates decision-making to units closest to the situation, granting latitude to deviate from detailed orders in service of higher-level intent [Shamir01102010]. Conceptually, each layer of management is an abstraction layer: when control is delegated, feedback loops shorten because fewer layers sit between the controlling entity and local conditions. This supports the broader hypothesis that delegation is a key ingredient of biological self-organisation that is largely absent from mainstream contemporary artificial intelligence. See ESM Note S7 for a concrete historical illustration (Nordbat 2).
@@ 3 What is Adaptation?
To make accurate inferences a system must be able to tell the difference between what it has caused, and what it has merely observed [merker2005, pearl2018, dawid2002]. It is often difficult to determine cause. ‘Confounding’ occurs when an event appears to be the cause of the effect we’re trying to predict, when in truth it is merely correlated. This can lead one to draw erroneous conclusions.
@@
By actively intervening in the environment, a system can experiment to establish cause [^21] Assuming such experimentation is enough to mitigate ambiguity and eliminate unknowns, and that a heuristic such as weakness is applied to ensure the system converges on variable definitions which represent cause accurately [bennett2023c].. Then, it can construct a causal graph. There are two ways to do that. One option would be to assume the nodes (variables, actions and other abstract objects and properties), and then learn the causal relations between them. This might work well in theory [richens2024robust], but it requires foreknowledge of objects and properties between which causal relations can be established. The alternative is to assume the causal relations (the edges in the graph), and then learn the variables. This needs only attraction and repulsion from physical states [bennett2023c, bennettwelshciaunica2024, bennett2025thesis], otherwise known as valence. I take this latter approach.
@@
Learning nodes to fit a graph is like figuring out which computer program [^22] Computation in this paper is about constraints, like how slime mold can be made to “compute” by imposing the right constraints on them[nakagaki2000]. I’m not assuming Turing machines. was used to generate a sequence. Assume I have fed a set of inputs into a computer. For each input the computer responded with a particular output. The computer runs a ‘policy’ program which maps inputs to outputs [^23] I call such programs policies, but this should not be confused with the more specific definition policy in reinforcement learning [sutton2018].. That policy is the cause of the input-output history observed so far. I want to know the policy that caused historical outputs so that I can use it to predict future outputs. The problem is that more than one policy could have produced the history we’ve observed, and each of those policies produce different outputs when I give the computer inputs it hasn’t encountered before. From a naive Bayesian perspective, all of them are equally likely to have caused the history I saw. If I wish to know which policy actually caused the outputs, I need a way to whittle down the remaining possibilities.
@@
There are two ways to compare policies that look the same with respect to historical data, but differ with respect to possible future data [bennett2024b]. One is to compare their form, and choose policies which are simpler in the sense of being more compressible. However, in an interactive setting simplicity is an unreliable indicator of plausibility [leike2015, bennett2024b]. The alternative is to compare policies by their function, in terms of how they behave given inputs for which I do not have historical data. Adaptive systems must satisfy functional constraints [sole2022]. I can compare systems by these constraints. Each policy implies a constraint on possible outputs given a set of possible inputs. The weaker that constraint is, the more variety of future behaviour it permits. Put another way, if two policies satisfy a goal, then the weaker policy satisfies it in more possible worlds [^24] To the extent that such possible worlds can be differentiated from one another within the subjective confines of an abstraction layer.. Previous research showed that if two policies are identical with respect to historical data, then the policy most likely to have generated historical data is the weaker [^25] Formal definitions and proofs for weakness and w-maxing are in [bennett2024appendices, bennett2025thesis]; an extended intuition is in ESM Supplementary Note S3. of the two [bennett2023c, bennett2025d, bennett2025e]. The policy most likely to have caused behaviour is the weakest [bennett2023c], and in experiments a system which preferred weaker policies outperformed one that preferred the simpler policies by $110-500\%$ [bennett2023b]. This has been used to establish an upper bound on embodied ‘intelligence’ [bennett2024a]. I will use the term ‘w-maxing’ to mean a system optimises for weak policies, and ‘simp-maxing’ to mean a system optimises for simpler policies [^26] Basically minimum description length policies [rissanen1978]., to conform to the terminology used in previous work [bennett2025b, bennett2025thesis] [^27] For formal definitions of learning and efficiency used in this framework, see [bennett2024appendices] and the thesis [bennett2025thesis].. Put another way, if two hypotheses appear equally good with respect to past data, the w-maxing system will use policy weakness as a heuristic to decide which policy to use, whereas a simp-maxing system will use simplicity.
@@
An abstraction layer can be understood as a border between form and function. For example a program takes a form when it is written in a language, but when it is run on a computer it has a particular function or effect upon the world [^28] For example, hardware like an Apple M2 system on a chip interprets software to physical effect, based on the ARM instruction set architecture.. Simp-maxing is a matter of form, while w-maxing is a matter of function [bennett2025b]. When we account for interpretation between form and function it becomes clear that software is a physical thing, like a tool. A tool can serve more than one purpose. For example, a hammer is both a weapon and a paperweight. This is what “weakness” measures. A tool that implies a weaker constraint can complete more tasks [bennettmaruyama2022a] than another tool [^29] To be exact, a weaker policy completes a superset of tasks, rather than just more tasks. This is because policies are compared in the context of an ostensive definition.. By preferring weaker policies, one avoids unnecessary commitments, meaning the weakest policies complete a ‘wider range of tasks’ than others [^30] See ESM Supplementary Note S3 for extended discussion of Bennett’s Razor, weakness, and related connections (e.g. maximum entropy).. To complete a ‘wider range of tasks’ in the embodied sense, a tool must be both more sample and more energy efficient, and so the weakest policy is the most efficient in both senses. Put another way, the same physical matter can simultaneously contribute to many different tasks or computations superimposed over each other at different scales. A ‘weaker’ policy serves more masters.
@@
ML systems do not typically optimise for either weaker constraints on function (w-maxing), or simplicity of form (simp-maxing). This may explain some of their relative inefficiency [bennett2022a, bennett2023b]. Even if they were w-maxing, the capacity for w-maxing is severely limited by the abstraction layer. An abstraction layer can be a “causal isolator” [friston2023] between a software “mind” at higher levels of abstraction, and its environment at lower levels. Embodiment by another name. Different abstraction layers permit different upper bounds on policy weakness, in the same way that two bodies might differ in how able they are to achieve a given goal. Different embodiments give different upper bounds on ‘intelligence’ [bennett2024a].
@@
That biological self-organising systems adapt efficiently means they are able to embody weaker policies than today’s ML systems. This capacity hinges on the abstraction layer [bennett2024a]. An abstraction layer is like a window through which only part of the world can be seen. It conveys an inductive bias that helps or hinders adaptation. When I optimise the system at a lower level of abstraction, this is like saying I can “move” this window to convey a different inductive bias. The bigger the window, the “weaker” the policy it is possible to learn. What this means is that intelligence is primarily a matter of the lens through which we interact with a problem. Hardware, not just software. In enactivist terms, an intelligent choice of abstraction layer is one that reduces the “big world” of all information to a “small world” of relevant information [vervaeke2012, vervaeke2013a, vervaeke2013b]. In a typical computer, lower levels of abstraction tend to be static. The window cannot move. I will show a static abstraction layer limits adaptability at higher levels of abstraction by limiting how weak the policies embodied by the system can be. Adaptability at high levels requires adaptability at low levels, and delegation of control facilitates this.
@@ 4 Formal Framework and Definitions
The formal definitions below are required for the core claim and for Theorem 5.1. The surrounding narrative provides intuition, but the theorem and its assumptions depend on these definitions. However this paper is based on a formal mathematical framework and it does contain a very simple proof supporting the claim. The formal details are also useful to connect this work here to related mathematical and experimental work [bennett2025thesis, bennett2023b, bennett2023c, bennett2024a, bennett2024b, bennettwelshciaunica2024, bennett2025e].
Stack Theory describes an adaptive system or computation by the binary constraints or boundary conditions that define it. It formalises embodiment as a choice of abstraction layer: a finite vocabulary of possible constraints (programs) representing the possible configurations of a body embedded in its environment [bennett2025thesis], which are used to express “statements” constraining the body. Policies and tasks are then defined using statements in this embodied language. Delegation, distribution and abstraction are captured by the task hierarchy. A longer conceptual motivation (including the environment/states/programs framing and links to enactivist and pancomputational interpretations) is provided in the Electronic Supplementary Materials (ESM) Note S5 and is not required to follow the theorem below.
For example, if a human raises their arm, then they are expressing a statement in their embodied language. That statement constrains what the body does next. The vocabulary $\mathfrak{v}$ of programs express this physical constraint. The inputs, outputs and policies of a system are all just subsets of its vocabulary. A $\mathfrak{v}$-task $\langle I,O\rangle$ is a set of inputs and corresponding outputs. If the $\mathfrak{v}$-task describes a ‘fit’ organism, then the outputs it contains are ‘correct’ according to natural selection. A system [ciaunica2023b] that conformed to the free energy principle (FEP) would be a $\mathfrak{v}$-task for which $\langle I,O\rangle$ minimised surprisal. In other words, a $\mathfrak{v}$-task can describe an organism as behaviour that serves homeostatic and reproductive goals.
For readers who wish to check the formal dependencies (see also [bennett2025thesis, bennett2024appendices]), the key definitions used in the theorem are summarised below.
@@
    A set $\Phi$ is assumed, whose elements are called states. A declarative program is a set $f$ of states, and $f$ is true relative to $\phi\in\Phi$ iff $\phi\in f$. $P=2^{\Phi}$ is the set of all programs. Embodiment is formalised as an abstraction layer, for which I single out a finite set $\mathfrak{v}\subseteq P$ called a vocabulary. A choice of vocabulary $\mathfrak{v}$ implies a formal language
    $L_{\mathfrak{v}}=\{l\subseteq\mathfrak{v}:\bigcap l\neq\emptyset\}.$
@@
    Members of $L_{\mathfrak{v}}$ are called statements, and a statement $l$ is considered true relative to a state $\phi$ iff $\phi\in\bigcap l$. A completion of a statement $l_{1}$ is any statement $l_{2}$ s.t. $l_{1}\subseteq l_{2}$.
    The extension of a statement $l$ is the set $E_{l}$ of all its completions, s.t. $x\in L_{\mathfrak{v}}$ has
    $E_{x}=\{y\in L_{\mathfrak{v}}:x\subseteq y\}.$
    The extension of a set $X\subseteq L_{\mathfrak{v}}$ of statements is
    $E_{X}=\bigcup\limits_{x\in X}E_{x}.$
    The weakness of a policy $x$ is the cardinality of its extension $\lvert E_{x}\rvert$.
@@
    A $\mathfrak{v}$-task $\alpha$ is a pair $\langle I_{\alpha},O_{\alpha}\rangle$ where $I_{\alpha}\subset L_{\mathfrak{v}}$ are the inputs, $E_{I_{\alpha}}$ is the set of outputs available given $I_{\alpha}$, and $O_{\alpha}\subset E_{I_{\alpha}}$ are the correct outputs. All tasks form a hierarchy. If $\alpha\sqsubset\omega$ then $\omega$ is a parent of $\alpha$, meaning $I_{\alpha}\subset I_{\omega}$ and $O_{\alpha}\subseteq O_{\omega}$. A parent is higher level than its children. A parent task is delegated to child tasks, and as each task is embodied this describes control being delegated within a system.
@@
    A policy is a statement $\pi$ and a policy is correct for $\alpha$ iff
    $\pi\in\Pi_{\alpha}=\{\pi\in L_{\mathfrak{v}}:E_{I_{\alpha}}\cap E_{\pi}=O_{\alpha}\}.$
    Given input $i\in I_{\alpha}$ and a policy $\pi$, to infer is to choose $o\in E_{i}\cap E_{\pi}$. A task is complete if $o\in O_{\alpha}$. Given a task definition $\alpha$ s.t. $\alpha\sqsubset\omega$, to learn is to choose a policy $\pi\in\Pi_{\alpha}$, and $\omega$ has been learned from $\alpha$ if one chooses $\pi\in\Pi_{\omega}$.
@@
  6. 6.
@@
  7. 7.
@@
  8. 8.
@@
  • The stack of layers of increasing abstraction is a sequence of uninstantiated tasks $\langle\lambda^{0},\lambda^{1},\dots,\lambda^{n}\rangle$ s.t. $\lambda^{i+1}\sqsubset\lambda^{i}$. Lower in the stack is a lower level of abstraction.
  • $\mathfrak{f}:2^{P}\times L_{\mathfrak{v}}\rightarrow 2^{P}$ is a function that takes a vocabulary $\mathfrak{v}$ and a policy $\pi\in L_{\mathfrak{v}}$, and returns a vocabulary
    $\mathfrak{v^{\prime}}=\{f\in P:\exists o\in E_{\pi}(\bigcap o=f)\}.$
  • The state of an MLA is a sequence of policies $\langle\pi^{0},\pi^{1},\dots,\pi^{n}\rangle$ and a sequence of vocabularies $\langle\mathfrak{v}^{0},\mathfrak{v}^{1},\dots,\mathfrak{v}^{n}\rangle$ such that
@@
I say the MLA is over-constrained when there exists $i<n$ such that $\Pi_{\lambda^{i}(\mathfrak{v}^{i})}=\emptyset$, and multilayer-causal-learning (MCL) is when the MLA is not over-constrained.

### v6

@@ 2.4 Reasons to Suspect Delegation Facilitates Adaptation
Theorem 5.1 provides the formal claim. Here I briefly motivate why delegation should matter. Because perceived complexity is defined by the abstraction layer through which a system interacts with its environment [17], measures of intelligence that ignore embodiment are incomplete; an “objective” measure must be enactive [86].
First, adaptation closer to the substrate can exploit more degrees of freedom. Cutting out an abstraction layer can yield large efficiency gains (C vs Python) [65]; reconfigurable substrates such as FPGAs push adaptation further down the stack [63].
@@
Second, delegation can improve robustness via graceful degradation, avoiding single points of failure [55]. For example, to control a drone swarm I could single out one drone to be the ‘leader’, and have it run a program that co-ordinates the other drones. However that would make the swarm vulnerable. If the leader drone were damaged, the swarm would de-cohere. In contrast if I can get the swarm to co-ordinate through entirely decentralised control, individual drones negotiating with those around them, then the swarm can remain coherent regardless of how many drones are disabled. This sort of decentralised swarm control can also be observed in nature, such as with flocks of starlings which fly in complex formations [25]. In other words, the swarm ‘adapts’ to adverse conditions.
@@
Third, in human organisations delegation shortens feedback loops. In military doctrine, “mission command” delegates decision-making to units closest to the situation, granting latitude to deviate from detailed orders in service of higher-level intent [79]. This trades central control for responsiveness. See ESM Note S7 for a concrete historical illustration (Nordbat 2 [47, 48]).
These are all intuition-building thought experiments rather than empirical proof, but they clarify the mechanism the theorem captures. To adapt at high levels, a system must retain enough slack at lower levels.
@@ 3 What is Adaptation?
To make accurate inferences a system must be able to tell the difference between what it has caused, and what it has merely observed [60, 72, 31]. It is often difficult to determine cause. ‘Confounding’ occurs when an event appears to be the cause of the effect we’re trying to predict, when in truth it is merely correlated. This can lead one to draw erroneous conclusions.
@@
By actively intervening in the environment, a system can experiment to establish cause [^3] Assuming such experimentation is enough to mitigate ambiguity and eliminate unknowns, and that a heuristic such as weakness is applied to ensure the system converges on variable definitions which represent cause accurately [14].. Then, it can construct a causal graph. There are two ways to do that. One option would be to assume the nodes (variables, actions and other abstract objects and properties), and then learn the causal relations between them. This might work well in theory [74], but it requires foreknowledge of objects and properties between which causal relations can be established. The alternative is to assume the causal relations (the edges in the graph), and then learn the variables [14, 19, 18]. This needs only attraction and repulsion from physical states [12, 19], otherwise known as valence. I take this latter approach. While causal discovery of this sort remains difficult to engineer in open-ended settings (e.g. unobserved confounders and combinatorial hypothesis spaces), I treat this as a modelling stance and focus on the stack-level delegation constraint formalised in Theorem 5.1.
@@
Learning nodes to fit a graph is like figuring out which computer program was used to generate a sequence. Assume I have fed a set of inputs into a computer. For each input the computer responded with a particular output. The computer runs a ‘policy’ program which maps inputs to outputs [^4] I call such programs policies, but this should not be confused with the more specific definition policy in reinforcement learning [84].. That policy is the cause of the input-output history observed so far. I want to know the policy that caused historical outputs so that I can use it to predict future outputs. The problem is that more than one policy could have produced the history we’ve observed, and each of those policies produce different outputs when I give the computer inputs it hasn’t encountered before. From a naive Bayesian perspective, all of them are equally likely to have caused the history I saw. If I wish to know which policy actually caused the outputs, I need a way to whittle down the remaining possibilities.
@@
There are two ways to compare policies that look the same with respect to historical data, but differ with respect to possible future data [17]. One is to compare their form, and choose policies which are simpler in the sense of being more compressible. However, in an interactive setting simplicity is an unreliable indicator of plausibility [56, 17]. The alternative is to compare policies by their function, in terms of how they behave given inputs for which I do not have historical data. Adaptive systems must satisfy functional constraints [83]. I can compare systems by these constraints. Each policy implies a constraint on possible outputs given a set of possible inputs. The weaker that constraint is, the more variety of future behaviour it permits. Put another way, if two policies satisfy a goal, then the weaker policy satisfies it in more possible worlds [^5] To the extent that such possible worlds can be differentiated from one another within the subjective constraint of an abstraction layer.. Previous research showed that if two policies are identical with respect to historical data, then the policy most likely to have generated historical data is the weaker [21, 19] (see ESM Supplementary Note S3 for formal details.) of the two [14, 20, 18]. The policy most likely to have caused behaviour is the weakest [14], and in experiments a system which preferred weaker policies outperformed one that preferred the simpler policies by $110-500\%$ [15]. This has been used to establish an upper bound on embodied ‘intelligence’ [16]. I will use the term ‘w-maxing’ to mean a system optimises for weak policies, and ‘simp-maxing’ to mean a system optimises for simpler (minimum description length [75]) policies, to conform to the terminology used in previous work [22, 19, 21]. Put another way, if two hypotheses appear equally good with respect to past data, the w-maxing system will use policy weakness as a heuristic to decide which policy to use, whereas a simp-maxing system will use simplicity.
Weakness is not the same thing as simplicity. Simplicity is a property of a policy’s description in a chosen encoding. Weakness is a property of the policy’s remaining completion space inside the embodied language $L_{\mathfrak{v}}$. Two correct policies can therefore have very different weakness even when their descriptions are equally short, or when one is syntactically longer. Weakness is also not mere truth set size. It counts how many distinct further commitments remain available to the agent given what it can express and implement at that abstraction layer. In this sense, w-maxing implements a least commitment heuristic stated at the level of function, rather than at the level of syntax. Since the completion structure depends on the vocabulary, weakness naturally incorporates embodiment. Changing sensors, effectors, or internal degrees of freedom changes $\mathfrak{v}$ and changes which refinements count as distinct. In the uniform case where all remaining completions are treated as equally plausible, maximising weakness coincides with a maximum entropy preference over viable futures.
@@
An abstraction layer can be understood as a border between form and function. For example a program takes a form when it is written in a language, but when it is run on a computer it has a particular function or effect upon the world. Simp-maxing is a matter of form, while w-maxing is a matter of function [22]. When we account for interpretation between form and function it becomes clear that software is a physical thing, like a tool. A tool can serve more than one purpose. For example, a hammer is both a weapon and a paperweight. This is what “weakness” measures. A tool that implies a weaker constraint can complete more tasks [10] than another tool. By preferring weaker policies, one avoids unnecessary commitments, meaning the weakest policies complete a ‘wider range of tasks’ than others (see ESM Note S3 for discussion of weakness and related connections (e.g. maximum entropy).). To complete a ‘wider range of tasks’ in the embodied sense, a tool must be both more sample and more energy efficient, and so the weakest policy is the most efficient in both senses. Put another way, the same physical matter can simultaneously contribute to many different tasks or computations superimposed over each other at different scales. A ‘weaker’ policy serves more masters.
@@
ML systems do not typically optimise for either weaker constraints on function (w-maxing), or simplicity of form (simp-maxing). This may explain some of their relative inefficiency [13, 15]. Even if they were w-maxing, the capacity for w-maxing is severely limited by the abstraction layer. An abstraction layer can be a “causal isolator” [35] between a software “mind” at higher levels of abstraction, and its environment at lower levels. Embodiment by another name. Different abstraction layers permit different upper bounds on policy weakness, in the same way that two bodies might differ in how able they are to achieve a given goal. Different embodiments give different upper bounds on ‘intelligence’ [16].
@@
That biological self-organising systems adapt efficiently means they are able to embody weaker policies than today’s ML systems. This capacity hinges on the abstraction layer [16]. An abstraction layer is like a window through which only part of the world can be seen. It conveys an inductive bias that helps or hinders adaptation. When I optimise the system at a lower level of abstraction, this is like saying I can “move” this window to convey a different inductive bias. The bigger the window, the “weaker” the policy it is possible to learn. What this means is that intelligence is primarily a matter of the lens through which we interact with a problem. Hardware, not just software. In enactivist terms, an intelligent choice of abstraction layer is one that reduces the “big world” of all information to a “small world” of relevant information [95, 94, 93]. In a typical computer, lower levels of abstraction tend to be static. The window cannot move. I will show a static abstraction layer limits adaptability at higher levels of abstraction by limiting how weak the policies embodied by the system can be. Adaptability at high levels requires adaptability at low levels, and delegation of control facilitates this.
@@ 4 Formal Framework and Definitions
The core claim and Theorem 5.1 depend upon the Stack Theory primitives defined below. The surrounding narrative provides intuition. The ESM contains additional motivation, proofs, and related technical material [19, 15, 14, 16, 17, 12, 18].
Stack Theory describes an adaptive system in terms of the binary constraints or boundary conditions that define it. Embodiment is a choice of abstraction layer. This is a finite vocabulary of possible constraints (programs) representing the possible configurations of a body embedded in its environment [19]. These programs are used to express “statements” constraining the body. Policies and tasks are defined as statements in this embodied language, and delegation is captured by how task constraints are arranged across layers. ESM Note S5 provides a longer conceptual motivation and links to related perspectives.
For example, if a human raises their arm, then they are expressing a statement in their embodied language. That statement constrains what the body does next. The vocabulary $\mathfrak{v}$ of programs express this physical constraint. The inputs, outputs and policies of a system are all just subsets of its vocabulary. A $\mathfrak{v}$-task $\langle I,O\rangle$ is a set of inputs and corresponding outputs. If the $\mathfrak{v}$-task describes a ‘fit’ organism, then the outputs it contains are ‘correct’ according to natural selection. A $\mathfrak{v}$-task $\langle I,O\rangle$ can represent biological selection pressures by specifying which outputs are viable for which inputs. For example, an FEP-style agent can be treated as a task whose admissible futures minimise surprisal under an appropriate generative model [27, 39].
For readers who wish to check the formal dependencies (see also [19, 21]), the key definitions used in the theorem are summarised below.
@@
    Environment, programs, vocabularies. Fix a nonempty set $\Phi$ of mutually exclusive states. A (declarative) program is a set $p\subseteq\Phi$, and $p$ is true at $\phi\in\Phi$ iff $\phi\in p$. Let $P:=2^{\Phi}$ be the set of all programs. Embodiment is formalised as an abstraction layer, for which I single out a finite vocabulary $\mathfrak{v}\subseteq P$.
    Intuition. A program is a set of states, and a vocabulary is the finite set of predicates the layer can express.
@@
    Statements and truth sets. The language generated by $\mathfrak{v}$ is
    $L_{\mathfrak{v}}:=\{\;l\subseteq\mathfrak{v}:T(l)\neq\emptyset\;\},\qquad T(l):=\bigcap_{p\in l}p\subseteq\Phi,$
    where $T(l)$ is the truth set (truth conditions) of the (conjunctive) statement $l$. By convention, $T(\emptyset)=\Phi$. A statement $l$ is true at state $\phi$ iff $\phi\in T(l)$.
    Intuition. A statement is a conjunction of vocabulary programs. Its truth set is the states where all of them hold.
@@
    Completions, extensions, weakness. A completion of a statement $l$ is any statement $l^{\prime}\in L_{\mathfrak{v}}$ such that $l\subseteq l^{\prime}$. The extension of $l$ is the set of all its completions.
    $\mathrm{Ext}(l):=\{\;l^{\prime}\in L_{\mathfrak{v}}:l\subseteq l^{\prime}\;\}.$
    For a set $X\subseteq L_{\mathfrak{v}}$, define $\mathrm{Ext}(X):=\bigcup_{x\in X}\mathrm{Ext}(x)$. (Throughout, $\mathrm{Ext}(\,\cdot\,)$ is always taken in the ambient language $L_{\mathfrak{v}}$ determined by context.) The weakness of a policy $\pi$ is the cardinality of its extension $|\mathrm{Ext}(\pi)|$.
    Intuition. A completion adds further constraints while staying compatible with what has already been fixed. Weakness counts how many distinct completions remain available in the agent’s own embodied language. It is therefore a measure of optionality or autonomy at that abstraction layer. Weakness is not a syntactic notion like description length. A policy can be very short yet highly committal, or long yet weak, depending on how the vocabulary factors the world. Changing embodiment changes the vocabulary and therefore changes which refinements count, which is intended.
@@
    Tasks and hierarchy. A $\mathfrak{v}$-task $\alpha$ is a pair $\langle I_{\alpha},O_{\alpha}\rangle$ where $I_{\alpha}\subseteq L_{\mathfrak{v}}$ is a set of admissible inputs and $O_{\alpha}\subseteq\mathrm{Ext}(I_{\alpha})$ is a set of correct outputs. Tasks form a hierarchy. If $\alpha\sqsubset\omega$ then $\omega$ is a parent of $\alpha$, meaning $I_{\alpha}\subseteq I_{\omega}$ and $O_{\alpha}\subseteq O_{\omega}$.
    Intuition. A task specifies which outputs count as correct for each input. The hierarchy orders tasks by adding constraints.
@@
    Policies, correctness, inference, learning. A policy is a statement $\pi\in L_{\mathfrak{v}}$. A policy is correct for $\alpha$ iff
    $\pi\in\Pi_{\alpha}:=\{\pi\in L_{\mathfrak{v}}:\mathrm{Ext}(I_{\alpha})\cap\mathrm{Ext}(\pi)=O_{\alpha}\}.$
    Given input $i\in I_{\alpha}$ and a policy $\pi$, to infer is to choose $o\in\mathrm{Ext}(i)\cap\mathrm{Ext}(\pi)$. A task is complete if $o\in O_{\alpha}$. Given $\alpha\sqsubset\omega$, to learn is to choose a policy $\pi\in\Pi_{\alpha}$, and $\omega$ has been learned from $\alpha$ if one chooses $\pi\in\Pi_{\omega}$.
    Intuition. A policy is a statement that filters which outputs are allowed. Learning chooses a policy whose completions match the observed input output pairs.
  6. 6.
@@
    Intuition. The point of this example is that tighter constraints at one layer limit how weak a compatible higher-layer policy can be. This is why delegation matters for efficiency.
  7. 7.
@@
    Intuition. Collective identity is the existence of at least one policy that is simultaneously viable for the parts and for the whole.
  8. 8.
@@
    Intuition. An uninstantiated task is a task template that can be applied to different vocabularies. This lets the same selection pressure be expressed across layers.
  9. 9.
@@
    Intuition. Utility increases when the correct-output set can be realised by many policies. This captures adaptability under the embodied language.
@@
  • The stack of layers of increasing abstraction is a sequence of uninstantiated tasks $\langle\lambda^{0},\lambda^{1},\dots,\lambda^{n}\rangle$ such that $\lambda^{i+1}\sqsubset\lambda^{i}$. Lower in the stack is a lower level of abstraction.
  • Given a vocabulary $\mathfrak{v}\subseteq P$ and a policy $\pi\in L_{\mathfrak{v}}$, the induced abstractor is
    $\mathfrak{f}(\mathfrak{v},\pi):=\{\,T(o):o\in\mathrm{Ext}(\pi)\,\}\subseteq P.$
    That is, the next-layer vocabulary consists of the truth sets of completions compatible with $\pi$.
  • The state of an MLA is a sequence of policies $\langle\pi^{0},\pi^{1},\dots,\pi^{n}\rangle$ and a sequence of vocabularies $\langle\mathfrak{v}^{0},\mathfrak{v}^{1},\dots,\mathfrak{v}^{n}\rangle$ such that for each $i<n$,
@@
Intuition. The stack records tasks at increasing abstraction, and each layer is constrained by the layer below. The abstractor tells you which higher-level predicates become available once a lower-level policy has fixed which completions are possible. The state records the current policy and induced vocabulary at each layer.
I say the MLA is over-constrained when there exists $i<n$ such that $\Pi_{\lambda^{i}(\mathfrak{v}^{i})}=\emptyset$, and multilayer-causal-learning (MCL) is when the MLA is not over-constrained. Over-constraint means some layer cannot realise any policy compatible with the higher-level task. The MCL is the boundary where each layer still has at least one viable policy, so collective identity is possible.